OpenStudy (anonymous):

Please help! Simplify: (4x^2-14x-30)/(x^2-1) ÷ (2x^2+5x+3)/(x^2-6x+5)

hero (hero):

Please use parentheses when posting the question to separate the numerator from the denominator.

OpenStudy (anonymous):

Sorry.

hero (hero):

That's a lot better... if you don't do that, then it leaves your expressions open to unnecessary interpretation.

OpenStudy (anonymous):

Had to fix 4x^2-14x-30)/(x^2-1)

OpenStudy (anonymous):

Sorry.

hero (hero):

So, now we have \[\frac{4x^2 - 14x - 30}{x^2} \div \frac{2x^2 + 5x + 3}{x^2 - 6x + 5}\] Which can be re-written as \[\frac{4x^2 - 14x - 30}{x^2} \times \frac{x^2 - 6x + 5 }{2x^2 + 5x + 3}\] The next thing to do is factor, which afterwards you get: \[\frac{2(2x + 3)(x -5)}{x^2} \times \frac{(x - 1)(x- 5) }{(x + 1)(2x + 3)}\]

hero (hero):

From here, you cancel factors of one

OpenStudy (anonymous):

How did you figure out the 2(2x +3) (x-5)?

hero (hero):

4x^2 - 14x - 30 Factor out 2 which is common to all 3 terms: 2(2x^2 - 7x - 15) Next factor x^2 - 7x - 15: 2(2x^2 - 10x + 3x - 15) 2(2x(x - 5) + 3(x - 5)) 2((x - 5)(2x + 3)) 2(x - 5)(2x + 3)

OpenStudy (anonymous):

Thanks.

hero (hero):

Are you sure you understood it?

OpenStudy (anonymous):

That part. I am still working on the problem.

hero (hero):

Suppose you were given another quadratic to factor. Do you think you could factor it in the same manner?

OpenStudy (anonymous):

Maybe.

OpenStudy (anonymous):

Why is it (x+1) (2x+3) and not (x+3) (x+2)?

hero (hero):

Multiply (x + 3)(x + 2) to find out

hero (hero):

Use the distributive rule when multiplying a(b + c) = ab + ac Similarly: (x + 3)(x + 2) = x(x + 3) + 2(x + 3) = x^2 + 3x + 2x + 6 = x^2 + 5x + 6

hero (hero):

\[2x^2 + 5x + 3 \ne x^2 + 5x + 6\]